# Bulk-Boundary Correspondence and Exceptional Points for a Dimerized Hatano-Nelson Model with Staggered Potentials

### v1

@@ VI.2 $\mathcal{PT}$-symmetric case
In brief, the entanglement entropy has a rich behaviour in the $\mathcal{PT}$ symmetric model and is capable of detecting exceptional points and phase transitions. In the $\mathcal{PT}$-unbroken topological phase there are two protected eigenvalues at $1/2$, contributing $2\log 2$ to the entanglement entropy.
@@
Figure 15: Entanglement entropy for the $\mathcal{PT}$ symmetric case with PBC and $u=1$ as a function of $W_{L}=W_{R}=W$ for two fixed values of $V_{L}=V_{R}=V$ as indicated in the labels. The lattice size is $N=120$. The dashed grid lines indicate the position of real exceptional points. In the insets, we zoom into regions where the exceptional points are very close to each other.
VII Conclusion
@@
In this paper we have analytically studied a paradigmatic one-dimensional, non-Hermitian lattice model. Our goal was to illustrate many of the unique features of non-Hermitian systems in a concrete model and to do so largely based on analytical instead of purely numerical results. We derived the eigensystems both for closed (periodic and anti-periodic) as well as open boundary conditions. Our analytical results show how the Bloch waves in the closed case change into localized eigenstates for open boundaries thus explicitly demonstrating the non-Hermitian skin effect. Another common feature of non-Hermitian systems are exceptional points where both eigenvalues and -vectors coalesce. In the considered model, we have shown that such points typically occur when the complex gap closes and that the location of exceptional points, in general, differs between the closed and the open case. Surprisingly, we found in the $\mathcal{PT}$-symmetric phase that exceptional points can even become dense in the thermodynamic limit, forming an entire “exceptional phase”.
@@
One of the central results of this work is an explicit demonstration of a proper bulk-boundary correspondence for a non-Hermitian system. This point has lead to some confusion in the literature. Here we have shown that the winding numbers for the periodic system determine the number of protected singular values which go to zero in the thermodynamic limit. This number also corresponds to the number of exact zero-energy edge modes for a semi-infinite chain. However, the eigenspectrum of the finite chain does, in general, not converge to that of the semi-infinite one. Thus, no bulk-boundary correspondence between winding numbers and the eigenspectrum of a finite system exists, in contrast to Hermitian systems. The exact zero-energy edge states in the semi-infinite non-Hermitian case are not exact eigenstates for a finite system. They become hidden which means that they correspond to metastable states which are, however, very long-lived and thus highly relevant for our understanding of non-Hermitian systems; see also Ref. [25] where this issue was discussed more generally. It is also important to note that even if zero-energy eigenstates for a finite system do exist, they are typically not topologically protected. For the studied model we demonstrated this explicitly by adding a small perturbation to the Hamiltonian and showing that the eigenenergies shift while the topologically protected singular values are stable.
@@
Non-trivial topology in the studied non-Hermitian system thus leads to two observable and experimentally relevant phenomena: (1) Extremely long-lived stable zero energy states, and (2) non-trivial entanglement in systems with a gap between occupied and unoccupied energy bands.
Appendix A Some details for the open boundary case
$P_{N}(z)=\det_{N}\left(\mathcal{T}_{N}-z\right).$ (A.1)
The roots $P_{N}(z)=0$ are the quasienergies of the Hamiltonian (II.1). The characteristic polynomial (A.1) for $\gamma=0$ is special as it satisfies the recurrence relations
$P_{2j}(z)$ $=$ $(-\text{i}u-z)P_{2j-1}(z)-V_{L}V_{R}P_{2j-2}(z),$
$P_{2j-1}(z)$ $=$ $(\text{i}u-z)P_{2j-2}(z)-W_{L}W_{R}P_{2j-3}(z)$ (A.2)
$P_{j}(z)$ $=$ $\left((z+\text{i}u)(z-\text{i}u)-(V_{L}V_{R}+W_{L}W_{R})\right)P_{j-2}(z)$ (A.3)
-- $V_{L}V_{R}W_{L}W_{R}P_{j-4}(z),$
where appropriate initial conditions must be chosen.
On the other hand, following [45], one notes that equations (II.6,II.7) can be iterated to the same fourth order difference equation,
$r_{j+2}+\frac{V_{R}W_{R}}{V_{L}W_{L}}r_{j-2}$
$=\left(-\frac{(\text{i}u+\epsilon)(\text{i}u-\epsilon)+(V_{L}V_{R}+W_{L}W_{R})}{V_{L}W_{L}}\right)r_{j}.$ (A.4)
The difference equation (A) can be compared to (A.3) leading to the solution (IV.10).
In addition, following [45], introduce
$T_{n}(\alpha,\beta)=\frac{\alpha^{n+1}-\beta^{n+1}}{\alpha-\beta}$ (A.5)
and the sequence
$a_{n+2}-(\alpha+\beta)a_{n+1}+\alpha\beta a_{n}=0.$ (A.6)
Then, it follows that
$a_{n}=T_{n-2}(\alpha,\beta)a_{2}-\alpha\beta T_{n-3}(\alpha,\beta)a_{1}$ (A.7)
is the solution of the difference equation (A.6). Now choosing
$\alpha+\beta=\frac{(\text{i}u-\epsilon)(-\text{i}u-\epsilon)-(V_{L}V_{R}+W_{L}W_{R})}{V_{L}W_{L}},$ (A.8)
$\alpha\beta=\frac{V_{R}W_{R}}{V_{L}W_{L}},$ (A.9)
it follows that
$r_{2j}=T_{j-2}(\alpha,\beta)r_{4}-\alpha\beta T_{j-3}(\alpha,\beta)r_{2},$ (A.10)
$r_{2j-1}=T_{j-2}(\alpha,\beta)r_{3}-\alpha\beta T_{j-3}(\alpha,\beta)r_{1},$ (A.11)
is a solution of (II.6,II.7,A). The boundary conditions determine,
$r_{4}$ $=$ $(\epsilon-\text{i}u)\left(2\frac{W_{R}}{V_{L}}+\alpha+\beta\right).$ (A.12)
where $r_{1}=V_{L}$ is an arbitrary choice. The conditions (A.8) give $\{\epsilon,\beta\}$ in terms of $\alpha$,
$\epsilon$ $=$ ±$\pm\sqrt{V_{L}V_{R}+W_{L}W_{R}+V_{L}W_{L}(\alpha+\beta)-u^{2}},$
$\beta$ $=$ $\frac{V_{R}W_{R}}{V_{L}W_{L}}\frac{1}{\alpha}.$ (A.13)
Using the parametrization
$\alpha=\frac{\sqrt{V_{R}}\sqrt{W_{R}}}{\sqrt{V_{L}}\sqrt{W_{L}}}\exp(\text{i}\theta),$ (A.14)
and then
$T_{n}\left(\alpha,\frac{V_{R}W_{R}}{V_{L}W_{L}}\frac{1}{\alpha}\right)=\frac{V_{R}^{\frac{n}{2}}W_{R}^{\frac{n}{2}}}{V_{L}^{\frac{n}{2}}W_{L}^{\frac{n}{2}}}\frac{\sin((n+1)\theta)}{\sin(\theta)}$ (A.15)
in (A.10,A.11) we find (IV.10,LABEL:vLodd) with the initial conditions (A).
The boundary condition at the end of the chain gives a condition on the parameter $\theta$, which depends on whether $N$ is odd or even. For odd $N$, the condition $x_{L+1}=0$ comes from (IV.10) with $12n=L+$ 1 and gives the quantized solution (IV.8). For even $N$, the condition $x_{L+1}=0$ comes from (LABEL:vLodd) with $12n-1=L+$ 1 and leads to the transcendental equation (IV.9). The solution (IV.12,LABEL:vLoddL) is obtained analogously.
Appendix B Discriminant of the characteristic polynomial
@@
$\text{disc}(P_{N}(z))=\text{disc}\left(\det_{N}\left(\mathcal{T}_{N}-z\right)\right)$ (B.1)
vanishes. The discriminant is given by the determinant of the Sylvester matrix [56], and it is in general an intricate expression of the parameters of the model. The vanishing of the discriminant (B.1) is a necessary condition for the existence of an exceptional point. Of course, the discriminant vanishes when the quasienergies collide. Additionally, for the model considered in this paper, the eigenvectors always coalesce when the quasienergies collide, recall the expressions (III.13,III.17,IV.10). Thus, the collision of quasienergies is also sufficient for the existence of an exceptional point.
In the main text, the conditions for gap closing/exceptional points are determined from the simple characterization of the eigenvalues by (III.7) and (IV.4). Here these results are verified by explicitly computing the discriminant (B.1) using Mathematica for different boundary conditions.
B.1 Closed boundary conditions
Let us consider $\gamma=1$ and $u=0$. For a lattice with $N=10$ sites the discriminant (B.1) is given by,
$\text{disc}(P_{10}^{\text{(PBC)}}(z))$ $=$ $10^{10}\left(W_{L}^{5}+V_{R}^{5}\right)\left(V_{L}^{5}+W_{R}^{5}\right)$ (B.2)
@@ B.1 Closed boundary conditions
The roots of the factors $\left(W_{L}^{5}+V_{R}^{5}\right)\left(V_{L}^{5}+W_{R}^{5}\right)$ are clearly given by (III.26). The extra factor comes from (III.29), which can be seen by writing (III.29) like,
$\frac{V_{L}W_{L}}{V_{R}W_{R}}=e^{\text{i}(k+k^{\prime})}=e^{\frac{2}{5}\text{i}\pi(\ell+\ell^{\prime})}$ (B.3)
with $\ell+\ell^{\prime}=1,\dots,5$. As mentioned in the main text, this last factor produce degenerate eigenvalues of $\mathcal{T}_{N}$ but associated with different eigenvectors, thus they are not exceptional points.
@@
×$\times\left(-V_{L}^{3}W_{R}+V_{L}^{2}W_{R}^{2}-V_{L}W_{R}^{3}+V_{L}^{4}+W_{R}^{4}\right).$
The factor $\left(V_{L}^{5}W_{L}^{5}-V_{R}^{5}W_{R}^{5}\right)$ is also present for the non-zero field case, and is solved by (B.3), however it does not lead to exceptional points. To analyze the factor $\left(\left(W_{L}+V_{R}\right)\left(V_{L}+W_{R}\right)-u^{2}\right)U(u)$ we recall (III.25) for every possible quantized $k$ (five for $N=10$). For example, for $k=2\pi$,
$H_{1}(2\pi)H_{2}(2\pi)-u^{2}=\left(W_{L}+V_{R}\right)\left(V_{L}+W_{R}\right)-u^{2}.$ (B.6)
@@
B.2 Open boundary conditions
Now let us consider $N=11$, $\gamma=0$ (OBC) and $u=0$. In this case the discriminant (B.1) is given by,
$\text{disc}(P_{11}^{\text{(OBC)}}(z))=48922361856V_{L}^{20}V_{R}^{20}W_{L}^{20}W_{R}^{20}f_{1}^{3}f_{2}^{3}f_{3}^{3}$ (B.8)
@@ B.2 Open boundary conditions
$f_{3}$ $=$ $V_{L}^{2}V_{R}^{2}+V_{L}V_{R}W_{L}W_{R}+W_{L}^{2}W_{R}^{2}.$ (B.9)
Of course, the discriminant (B.8) vanishes when any of the couplings vanishes, consistent with our findings in the main text. The remaining factors can be verified by rewriting
$f_{1}f_{2}f_{3}=V_{L}^{5}V_{R}^{5}\sum_{j=0}^{5}\left(\frac{1}{\sqrt{\delta}}\right)^{2j}$ (B.10)
@@
$\text{disc}(P_{11}^{\text{(OBC)}}(z))$ $=$ $-48922361856V_{L}^{20}V_{R}^{20}W_{L}^{20}W_{R}^{20}$ (B.11)
@@
where $f_{j}$ is given by (B.2) and $g_{j}$ are given by
@@
$\prod_{m=1}^{5}u^{2}-\tilde{H}_{1}(\theta_{m})\tilde{H}_{2}(\theta_{m})=g_{1}g_{2}g_{3}$ (B.13)
@@
$\text{disc}(P_{12}^{\text{(OBC)}}(z))=4096V_{L}^{20}V_{R}^{20}W_{L}^{30}W_{R}^{30}U_{0}U_{1}(u)$ (B.14)
@@
Again, it is clear that the vanishing of any of the coupling parameters leads to a null discriminant (B.14). The factor $U_{0}$ can be written as
@@
$+\frac{34709}{153664}\frac{1}{\sqrt{\delta}^{6}}+\frac{279}{2401}\frac{1}{\sqrt{\delta}^{8}}+\frac{108}{2041}\frac{1}{\sqrt{\delta}^{10}}\Big{)}.$ (B.17)
@@
$+\frac{\bar{u}(\bar{u}-1)(3\bar{u}^{2}+\bar{u}-1)}{\sqrt{\delta}^{4}}-\frac{\bar{u}(\bar{u}+1)^{2}(\bar{u}-1)^{3}}{\sqrt{\delta}^{6}}\Bigg{)}.$ (B.18)
@@
$05$ 0.209455 i $-0.628308+0.408561\text{i}$
$0.580505+0.209455\text{i}$ $-0.628308-0.408561\text{i}$
$1.07762-0.24362\text{i}$ $-0.344178+0.638086\text{i}$
$1.07762+0.24362\text{i}$ $-0.344178-0.638086\text{i}$
@@
Table 2: We consider $\bar{u}=0.5$ First six lines, $+$ solutions, last six lines -- solutions.
@@
$0.642557-0.0955613\text{i}$ $-0.673088+0.294372\text{i}$
$0.642557+0.0955613\text{i}$ $-0.673088-0.294372\text{i}$
[…]
$1.5708+0.191123\text{i}$ $-0.694828\text{i}$
@@
$0.343297$ $-1.31139$
$1.10745-0.172812\text{i}$ $-0.36085+0.604218\text{i}$
$1.10745+0.172812\text{i}$ $-0.36085-0.604218\text{i}$

### v3

@@ VI.2 $\mathcal{PT}$-symmetric case
In brief, the entanglement entropy has a rich behavior in the $\mathcal{PT}$ symmetric model and is capable of detecting both exceptional points and phase transitions. In the $\mathcal{PT}$-unbroken topological phase there are two protected eigenvalues at $1/2$, contributing $2\log 2$ to the entanglement entropy.
@@
Figure 15: Entanglement entropy for the $\mathcal{PT}$ symmetric case with PBC and $u=1$ as a function of $W_{L}=W_{R}=W$ for two fixed values of $V_{L}=V_{R}=V$ as indicated in the labels. The lattice size is $N=120$. The dashed grid lines indicate the position of real exceptional points. In the insets, we zoom into regions where the exceptional points are very close to each other. In the thermodynamic limit, the exceptional points become close and $S_{N/2}$ is nowhere continuous.
VII Conclusions
In this paper we have analytically studied a paradigmatic one-dimensional, non-Hermitian lattice model. Our goal was to illustrate many of the unique features of non-Hermitian systems—in particular, the proper bulk-boundary correspondence and the exceptional points—in a concrete model and to do so largely based on analytical instead of purely numerical results. A particular focus of our work was to clarify the recently described relation between topological winding numbers for periodic systems and topologically protected singular values as well as hidden zero modes in open systems [22]. We have shown explicitly that the latter only become exact zero-energy eigenmodes in the limit of a semi-infinite chain. For a finite chain, they correspond to metastable states instead which is a crucial distinction for experiments.
@@ VII Conclusions
We derived the eigensystems both for closed (periodic and anti-periodic) as well as open boundary conditions. Our analytical results show how the Bloch waves in the closed case change into localized eigenstates for open boundaries thus explicitly demonstrating the non-Hermitian skin effect. Another common feature of non-Hermitian systems are exceptional points where both eigenvalues and -vectors coalesce. In the considered model, we have shown that such points typically occur when the complex gap closes and that the location of exceptional points, in general, differs between the closed and the open case. Surprisingly, we found in the $\mathcal{PT}$-symmetric case that exceptional points can even become dense in the thermodynamic limit, forming an entire novel “exceptional phase”. This phase has also other unique properties which are impossible in a Hermitian phase. We found, in particular, that transport is hyper-ballistic with a dynamical critical exponent $z=1/2$.
@@
As already mentioned above, one of the central results of this work is an explicit demonstration of a proper bulk-boundary correspondence for a non-Hermitian system. This point has lead to some confusion in the literature. Here we have shown that the winding numbers for the periodic system determine the number of protected singular values which go to zero in the thermodynamic limit. This number also corresponds to the number of exact zero-energy edge modes for a semi-infinite chain. However, the eigenspectrum of the finite chain does, in general, not converge to that of the semi-infinite one. Thus, no bulk-boundary correspondence between winding numbers and the eigenspectrum of a finite system exists, in contrast to Hermitian systems. The exact zero-energy edge states in the semi-infinite non-Hermitian case are not exact eigenstates for a finite system. They become hidden which means that they correspond to metastable states which are, however, very long-lived and thus highly relevant for our understanding of non-Hermitian systems; see also Ref. [22] where this issue was discussed more generally. It is also important to note that even if zero-energy eigenstates for a finite system do exist, they are typically not topologically protected. For the studied model we demonstrated this explicitly by adding a small perturbation to the Hamiltonian and showing that the eigenenergies shift while the topologically protected singular values are stable.
@@
A non-trivial topology in non-Hermitian systems thus leads to two observable and experimentally relevant phenomena: (1) Extremely long-lived metastable zero energy states, and (2) non-trivial entanglement in systems with a gap between occupied and unoccupied energy bands.
We end this section by mentioning some possible directions for future studies. First, it would be worthwhile to study the loci of exceptional points for other families of non-Hermitian free-fermionic models, including free fermions and parafermions in disguise [64, 65]. More importantly, it would be very interesting to extend these studies to the loci of exceptional points for non-Hermitian interacting models, see e.g. Ref. [66] for inspiration. Second, it would be important to go back to other models which have been studied in the literature using incorrect ideas of topology and to work out the proper bulk-boundary relations. For one specific example this has already been done in Ref. [22]. Finally, the scaling behavior of the entanglement entropy at the various critical lines observed in this work certainly deserves further investigations. The exceptional phase in the $\cal{PT}$-symmetric case is particularly intriguing, and studying this novel phase further might advance our understanding of critical behavior in non-Hermitian systems.
Acknowledgments
R.A.P. acknowledges support from the Rio Grande do Sul State Research Support Foundation (FAPERGS) and the National Council for Scientific and Technological Development (CNPq). J.S. acknowledges support by NSERC via the Discovery grants program.
Appendix A Discriminant of the characteristic polynomial
@@
$\text{disc}(P_{N}(z))=\text{disc}\left(\det_{N}\left(\mathcal{T}_{N}-z\right)\right)$ (A.1)
vanishes. The discriminant is given by the determinant of the Sylvester matrix [67], and it is in general an intricate expression of the parameters of the model. The vanishing of the discriminant (A.1) is a necessary condition for the existence of an exceptional point. Of course, the discriminant vanishes when the quasienergies collide. Additionally, for the model considered in this paper, the eigenvectors always coalesce when the quasienergies collide, recall the expressions (III.13,III.17,IV.10). Thus, the collision of quasienergies is also sufficient for the existence of an exceptional point.
In the main text, the conditions for gap closing/exceptional points are determined from the simple characterization of the eigenvalues by (III.7) and (IV.4). Here these results are verified by explicitly computing the discriminant (A.1) using Mathematica for different boundary conditions.
A.1 Closed boundary conditions
Let us consider $\gamma=1$ and $u=0$. For a lattice with $N=10$ sites the discriminant (A.1) is given by,
$\text{disc}(P_{10}^{\text{(PBC)}}(z))$ $=$ $10^{10}\left(W_{L}^{5}+V_{R}^{5}\right)\left(V_{L}^{5}+W_{R}^{5}\right)$ (A.2)
@@ A.1 Closed boundary conditions
The roots of the factors $\left(W_{L}^{5}+V_{R}^{5}\right)\left(V_{L}^{5}+W_{R}^{5}\right)$ are clearly given by (III.26). The extra factor comes from another type of degeneracy in the spectrum which can occur within each band $\epsilon^{\pm}$ for different quasimomentum $k$. Namely, one can impose
$H_{1}(k)H_{2}(k)=H_{1}(k^{\prime})H_{2}(k^{\prime})$ (A.3)
which gives another constraint on the model parameters. Note that these degeneracies do not lead to exceptional points, as the eigenvectors do not coalesce, recall (III.13,III.17). The extra factor in the discriminant is indeed given by (A.3) which vanishes for
$\frac{V_{L}W_{L}}{V_{R}W_{R}}=e^{\text{i}(k+k^{\prime})}=e^{\frac{2}{5}\text{i}\pi(\ell+\ell^{\prime})}$ (A.4)
with $\ell+\ell^{\prime}=1,\dots,5$.
@@
$\times\left(-V_{L}^{3}W_{R}+V_{L}^{2}W_{R}^{2}-V_{L}W_{R}^{3}+V_{L}^{4}+W_{R}^{4}\right).$
The factor $\left(V_{L}^{5}W_{L}^{5}-V_{R}^{5}W_{R}^{5}\right)$ is also present for the non-zero field case, and is solved by (A.4), however it does not lead to exceptional points. To analyze the factor $\left(\left(W_{L}+V_{R}\right)\left(V_{L}+W_{R}\right)-u^{2}\right)U(u)$ we recall (III.25) for every possible quantized $k$ (five for $N=10$). For example, for $k=2\pi$,
$H_{1}(2\pi)H_{2}(2\pi)-u^{2}=\left(W_{L}+V_{R}\right)\left(V_{L}+W_{R}\right)-u^{2}.$ (A.7)
@@
A.2 Open boundary conditions
Now let us consider $N=11$, $\gamma=0$ (OBC) and $u=0$. In this case the discriminant (A.1) is given by,
$\text{disc}(P_{11}^{\text{(OBC)}}(z))=48922361856V_{L}^{20}V_{R}^{20}W_{L}^{20}W_{R}^{20}f_{1}^{3}f_{2}^{3}f_{3}^{3}$ (A.9)
@@ A.2 Open boundary conditions
$f_{3}$ $=$ $V_{L}^{2}V_{R}^{2}+V_{L}V_{R}W_{L}W_{R}+W_{L}^{2}W_{R}^{2}.$ (A.10)
Of course, the discriminant (A.9) vanishes when any of the couplings vanishes, consistent with our findings in the main text. The remaining factors can be verified by rewriting
$f_{1}f_{2}f_{3}=V_{L}^{5}V_{R}^{5}\sum_{j=0}^{5}\left(\frac{1}{\sqrt{\delta}}\right)^{2j}$ (A.11)
@@
$\text{disc}(P_{11}^{\text{(OBC)}}(z))$ $=$ $-48922361856V_{L}^{20}V_{R}^{20}W_{L}^{20}W_{R}^{20}$ (A.12)
@@
where $f_{j}$ is given by (A.2) and $g_{j}$ are given by
@@
$\prod_{m=1}^{5}u^{2}-\tilde{H}_{1}(\theta_{m})\tilde{H}_{2}(\theta_{m})=g_{1}g_{2}g_{3}$ (A.14)
@@
$\text{disc}(P_{12}^{\text{(OBC)}}(z))=4096V_{L}^{20}V_{R}^{20}W_{L}^{30}W_{R}^{30}U_{0}U_{1}(u)$ (A.15)
@@
Again, it is clear that the vanishing of any of the coupling parameters leads to a null discriminant (A.15). The factor $U_{0}$ can be written as
@@
$+\frac{34709}{153664}\frac{1}{\sqrt{\delta}^{6}}+\frac{279}{2401}\frac{1}{\sqrt{\delta}^{8}}+\frac{108}{2041}\frac{1}{\sqrt{\delta}^{10}}\Big).$ (A.18)
@@
$+\frac{\bar{u}(\bar{u}-1)(3\bar{u}^{2}+\bar{u}-1)}{\sqrt{\delta}^{4}}-\frac{\bar{u}(\bar{u}+1)^{2}(\bar{u}-1)^{3}}{\sqrt{\delta}^{6}}\Bigg).$ (A.19)
@@
$0.580505-0.209455\text{i}$ $0.628308+0.408561\text{i}$
$0.580505+0.209455\text{i}$ $0.628308-0.408561\text{i}$
$1.07762-0.24362\text{i}$ $0.344178+0.638086\text{i}$
$1.07762+0.24362\text{i}$ $0.344178-0.638086\text{i}$
@@
Table 2: We consider $\bar{u}=0.5$ First six lines, ++ solutions, last six lines - solutions.
@@
$0.642557-0.0955613\text{i}$ $0.673088+0.294372\text{i}$
$0.642557+0.0955613\text{i}$ $0.673088-0.294372\text{i}$
$1.5708-0.191123\text{i}$ $0.694828\text{i}$
$1.5708+0.191123\text{i}$ $0.694828\text{i}$
@@
$0.343297$ 1.31139-1.31139
$1.10745-0.172812\text{i}$ $0.36085+0.604218\text{i}$
$1.10745+0.172812\text{i}$ $0.36085-0.604218\text{i}$
@@
Appendix B Some details for the open boundary case
$P_{N}(z)=\det_{N}\left(\mathcal{T}_{N}-z\right).$ (B.1)
The roots $P_{N}(z)=0$ are the quasienergies of the Hamiltonian (II.1). The characteristic polynomial (B.1) for $\gamma=0$ is special as it satisfies the recurrence relations
$P_{2j}(z)$ $=$ $(-\text{i}u-z)P_{2j-1}(z)-V_{L}V_{R}P_{2j-2}(z),$
$P_{2j-1}(z)$ $=$ $(\text{i}u-z)P_{2j-2}(z)-W_{L}W_{R}P_{2j-3}(z)$ (B.2)
$P_{j}(z)$ $=$ $\left((z+\text{i}u)(z-\text{i}u)-(V_{L}V_{R}+W_{L}W_{R})\right)P_{j-2}(z)$ (B.3)
- $V_{L}V_{R}W_{L}W_{R}P_{j-4}(z),$
where appropriate initial conditions must be chosen.
On the other hand, following [50], one notes that equations (II.7,II.8) can be iterated to the same fourth order difference equation,
$r_{j+2}+\frac{V_{R}W_{R}}{V_{L}W_{L}}r_{j-2}$
$=\left(-\frac{(\text{i}u+\epsilon)(\text{i}u-\epsilon)+(V_{L}V_{R}+W_{L}W_{R})}{V_{L}W_{L}}\right)r_{j}.$ (B.4)
The difference equation (B) can be compared to (B.3) leading to the solution (IV.10).
In addition, following [50], introduce
$T_{n}(\alpha,\beta)=\frac{\alpha^{n+1}-\beta^{n+1}}{\alpha-\beta}$ (B.5)
and the sequence
$a_{n+2}-(\alpha+\beta)a_{n+1}+\alpha\beta a_{n}=0.$ (B.6)
Then, it follows that
$a_{n}=T_{n-2}(\alpha,\beta)a_{2}-\alpha\beta T_{n-3}(\alpha,\beta)a_{1}$ (B.7)
is the solution of the difference equation (B.6). Now choosing
$\alpha+\beta=\frac{(\text{i}u-\epsilon)(-\text{i}u-\epsilon)-(V_{L}V_{R}+W_{L}W_{R})}{V_{L}W_{L}},$ (B.8)
$\alpha\beta=\frac{V_{R}W_{R}}{V_{L}W_{L}},$ (B.9)
it follows that
$r_{2j}=T_{j-2}(\alpha,\beta)r_{4}-\alpha\beta T_{j-3}(\alpha,\beta)r_{2},$ (B.10)
$r_{2j-1}=T_{j-2}(\alpha,\beta)r_{3}-\alpha\beta T_{j-3}(\alpha,\beta)r_{1},$ (B.11)
is a solution of (II.7,II.8,B). The boundary conditions determine,
$r_{4}$ $=$ $(\epsilon-\text{i}u)\left(2\frac{W_{R}}{V_{L}}+\alpha+\beta\right).$ (B.12)
where $r_{1}=V_{L}$ is an arbitrary choice. The conditions (B.8) give $\{\epsilon,\beta\}$ in terms of $\alpha$,
$\epsilon$ $=$ $\pm\sqrt{V_{L}V_{R}+W_{L}W_{R}+V_{L}W_{L}(\alpha+\beta)-u^{2}},$
$\beta$ $=$ $\frac{V_{R}W_{R}}{V_{L}W_{L}}\frac{1}{\alpha}.$ (B.13)
Using the parametrization
$\alpha=\frac{\sqrt{V_{R}}\sqrt{W_{R}}}{\sqrt{V_{L}}\sqrt{W_{L}}}\exp(\text{i}\theta),$ (B.14)
and then
$T_{n}\left(\alpha,\frac{V_{R}W_{R}}{V_{L}W_{L}}\frac{1}{\alpha}\right)=\frac{V_{R}^{\frac{n}{2}}W_{R}^{\frac{n}{2}}}{V_{L}^{\frac{n}{2}}W_{L}^{\frac{n}{2}}}\frac{\sin((n+1)\theta)}{\sin(\theta)}$ (B.15)
in (B.10,B.11) we find (IV.10,LABEL:vLodd) with the initial conditions (B).
The boundary condition at the end of the chain gives a condition on the parameter $\theta$, which depends on whether $N$ is odd or even. For odd $N$, the condition $x_{L+1}=0$ comes from (IV.10) with $2n=L+1$ and gives the quantized solution (IV.8). For even $N$, the condition $x_{L+1}=0$ comes from (LABEL:vLodd) with 2n1=L+12n-1=L+1 and leads to the transcendental equation (IV.9). The solution (IV.12,LABEL:vLoddL) is obtained analogously.